PLEASE NOTE! I ONLY NEED ANSWERS TO QUESTION #8

Kate and Claire, recent college graduates, are unable to find suitable jobs in their field of accounting. However, each has been involved with a small business of their own for the last several years, and have been doing very well. Kate is a talented seamstress, and has designed a line of fashionable blazers that are selling for $500/each. Kate remains shocked at how fast the orders are coming in, and wonders if this could be something big.   Claire also has a small but growing business. She manufactures synthetic leather belts that are selling for $50/each, and is similarly experiencing strong consumer interest. A few large retailers have started to place orders with both girls, and both are struggling to keep up with demand. Kate and Claire are wondering if they should combine their lines and start building the business together, since there is a high amount of overlap among their customers, and they could likely achieve some synergies by combining their marketing and customer service efforts. The belts go very well with the blazers. With a solid knowledge of their accounting basics, both have kept very thorough cost and marketing data. So they decided to pull it all together and analyze it.

Kate’s blazers – manufacturing data

# of Blazers

Total Manufacturing Costs

2019

400

$140,000

2018

350

130,000

2017

310

122,000

2016

240

108,000

2015

275

115,000

2014

250

106,000

Kate’s blazers – marketing data

# of Blazers

Total Marketing Costs

2019

400

$60,000

2018

350

55,000

2017

310

51,000

2016

240

44,000

2015

275

47,500

2014

250

45,000

Claire’s belts – manufacturing data

# of Belts

Total Manufacturing Costs

2019

1,700

$66,500

2018

1,400

56,000

2017

1,100

45,500

2016

1,000

42,000

2015

1,200

49,000

2014

900

38,500

Claire’s belts – marketing data

# of Belts

Total Marketing Costs

2019

1,700

$11,500

2018

1,400

10,000

2017

1,100

8,500

2016

1,000

8,000

2015

1,200

9,000

2014

900

7,500

DELIVERABLE:

Prepare a comprehensive report with the answers to each of the following questions. Provide all calculations in a well-organized manner, with the final answer for each part clearly stated in a complete sentence.

Total report should be between 2-3 pages.

QUESTIONS:

1. High-low cost estimation method

  1. Use the high-low method to estimate the per-unit variable costs and total fixed costs for the blazers.
  2. Use the high-low method to estimate the per-unit variable costs and total fixed costs for the belts.

NOTE: The variable and fixed costs for each product line, blazers and belts, contain both a manufacturing and a marketing component. However, the data is recorded separately, which means FOUR separate high-low analyses must be conducted.

2. Cost-Volume-Profit (CVP) analysis, single-product

  1. Use CVP analysis to calculate the break-even point in units for the blazers.
  2. Use CVP analysis to calculate the break-even point in units for the belts.

3. CVP, multiple-product setting

Merging the data together, it appears the sales mix is approximately 300 blazers and 1,200 belts each year.

For this CVP analysis, assume an additional $30,000 of combined fixed costs, this will be largely customer service costs.

  1. Calculate the break-even point for both product lines combined.

4. Cost classification

  1. Classify the manufacturing costs, marketing costs, and customer service costs either as production expenses or period expenses.
  2. For the period expenses, further classify them into either selling expenses or general and administrative expenses.

5. Sensitivity CVP analysis and production versus period expenses – Multiple-Product Setting

  1. If both variable and fixed production expenses (refer to Question #1) associated with the blazer increased by 5% (beyond the estimate from the high-low analysis), how many blazers and belts would need to be sold in order to earn a target income of $96,000? Assume the same sales mix and additional fixed costs in Question #3.

6. Margin of Safety

  1. Calculate the margin of safety (both in units and in sales dollars), assuming they sell 700 blazers and 2,500 belts next year. Ignore the suggested change in expenses in Question #5.
  2. What does this margin of safety mean?

7. Income statement

  1. Prepare a forecasted income statement (to the operating income line) for Kate and Claire, assuming they sell 700 blazers and 2,500 belts, and assuming the 5% cost increase for the blazer (variable and fixed production costs) and the additional $30,000 combined fixed costs.
  2. How would you suggest Kate and Claire divide the operating profit?

8. Degree of Operating Leverage

  1. Calculate the degree of operating leverage, with the same assumptions as Question #7.
  2. What does the DOL tell Kate and Claire? Do you think they should consider promotional events? Why or why not?
Answer #1

Solution 8 :-

a) As the question specifies to use high low cost estimation

For blazers (manufacturing data)

Variable cost of blazer = (highest activity cost - lowest activity cost) / (highest activity units - lowest activity units)

= (140000-108000)/(400-240)

= 32000/160

=200 per unit

Fixed cost of blazer= highest activity cost - (variable cost* highest activity units)

= 140000-(200*400)

= 140000-80000 = 60000

Blazer ( marketing cost)

Variable cost = (60000-44000)/(400-240)

= 16000/160

=100 per unit

Fixed cost = 60000-(100*400)

= 20000

For belts ( manufacturing cost)

Variable cost= (66500-38500)/(1700-900)

= 28000/800

= 35 per unit

Fixed cost = 66500-(1700*35)= 7000

For belts ( marketing cost)

Variable cost = (11500-7500)/(1700-900)

= 5 per unit

Fixed cost = 11500- ( 1700*5)

= 3000

Revised rate of blazer ( 5% increase in variable and fixed production cost provided)

Variable cost = 200+5% *200 = 210 per unit

Fixed cost. = 60000+5%*60000 = 63000

Degree of operating leverage = sales - variable cost/sales - variable cost - fixed cost

Total variable cost ( for blazer) = VC for manufacturing+ VC for marketing

= 210+100 = 310 per unit

Total fixed cost ( for blazer) = FC for manufacturing+FC for marketing

= 63000+ 20000= 83000

Total variable cost (for belt) = VC for manufacturing+VC for marketing

= 35+5 = 40 per unit

Total fixed cost for belt = FC for manufacturing+FC for marketing

= 7000+3000= 10000

Degree of operating leverage = (sales - variable cost)/sales - variable cost - fixed cost)

Sales = 700*500+ 2500*50= 350000+125000= 475000

Variable cost = 700*310 + 2500*40 = 217000+100000= 317000

Fixed cost = Total FC for belts + total FC for blazer+ additional FC

= 10000+83000+30000= 123000

DOL =( 475000-317000)/(475000-317000-123000)

=158000/35000

= 4.514

b) DOL of the combination tells that for increase in sales by 1% will increase the contribution by 4.51 %

They promotional events should be considered as the company is new and Kate and Claire are in the growing stage. Further as there is no direct information provided that provides the direct increase in sales data because of the marketing expenses it is difficult to quantify the profit emerging from using the promotional events.